OpenStudy (anonymous):

Solve the linear system using the Gauss-Jordan elimination method. -0.3x - 0.8y= -9 0.6x -1.4y= -15 The solution of the system is x= ___, y=____

OpenStudy (stamp):

- 0.3x - 0.8y = -9 0.6x - 1.4y = -15 2(-0.3x - 0.8y = -9) 0.6x - 1.4y = -15 -0.6x - 1.6y = -18 0.6x - 1.4y = -15 Add the two systems 0x - 3.0y = -33 -3.0y = -33 y = -33 / -3.0 y = 11 Now substitute y in for x in either of your equations in the system and solve for x.

OpenStudy (anonymous):

What is the system for x?

OpenStudy (anonymous):

@stamp

OpenStudy (anonymous):